A parallel-plate capacitor has capacitance 1.00 μF.

(a) How much energy is stored in the capacitor if it is connected to a 7.00-V battery?
μJ

(b) If the battery is disconnected and the distance between the charged plates doubled, what is the energy stored?
μJ

(c) The battery is subsequently reattached to the capacitor, but the plate separation remains as in part (b). How much energy is stored?
μJ
Answer #1

here,

capacitance , C = 1 uF = 1 * 10^-6 F

a)

potential difference , V = 7 V

the energy stored , U = 0.5 * C * V^2

U = 0.5 * 1 * 7^2 uJ

U = 24.5 uJ

b)

when the battery is disconnected , the charge stored remains the same

C = area * e0 /d

when the sepration between the plates is doubled

the new capacitance , C' = C /2

the energy stored , U' = 0.5 * Q^2 /C'

U' = 0.5 * (C * V)^2 /(C/2)

U' = 0.5 * 2 * 1 * 7^2 uJ

U' = 49 uJ

c)

C = area * e0 /d

when the sepration between the plates is doubled

the new capacitance , C' = C /2

when the battery remains connected, the potential difference remains the same

the energy stored , U' = 0.5 * C' * V^2

U' = 0.5 * ( 1/2) * 7^2 uJ

U' = 12.25 uJ